OpenStudy (anonymous):

Write the standard form of the equation of the line that is parallel to the given line, y = −2x + 6, and passes through 0,−2

OpenStudy (texaschic101):

you need to find the slope. In y = mx + b form, the m position stands for the slope. So what is your slope in y = -2x + 6 ?

OpenStudy (anonymous):

so what's the answer sir? :)

OpenStudy (texaschic101):

Since you are looking for a parallel line, the slope will be the same. you will then use : y - y1 = m(x - x1) slope(m) = -2 (0,-2) x1 = 0 and y1 = -2 now lets sub. y - (-2) = -2(x - 0) y + 2 = -2x + 0 y = -2x + 0 - 2 y = -2x - 2 standard form... y = -2x - 2 2x + y = -2 (standard form)

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

Done it perfectly sir, Thanks a lot :)

OpenStudy (anonymous):

yes, i have many questions sir :)

OpenStudy (texaschic101):

ask away ?

OpenStudy (anonymous):

Write the equation of the line, in slope-intercept form, that passes through the point −5,5 Ê Ë ˆ ¯ and has slope −3.

OpenStudy (texaschic101):

slope(m) = -3 (-5,5) x1 = -5 and y1 = 5 y - y1 = m(x - x1) y - 5 = -3(x - (-5) y - 5 = -3(x + 5) y - 5 = -3x - 15 y = -3x - 15 + 5 y = -3x - 10 <-- slope intercept form

OpenStudy (texaschic101):

need anymore help ?

OpenStudy (anonymous):

Yes madam wait :) ill post

OpenStudy (anonymous):

5. Graph the linear system and determine how many solutions it has: no solution, exactly one solution or infinitely many solutions. (3 points) a. x + y = 2 3x − 2y = 6 b. 3x + y = 1 2y = 2 − 6x c. 4x − 2y = 8 2x − y = 7

OpenStudy (texaschic101):

(a) x + y = 2 --> x = -y + 2 now sub -y + 2 in for x in the other equation 3x - 2y = 6 3(-y + 2) - 2y = 6 -3y + 6 - 2y = 6 -5y = 6 - 6 -5y = 0 y = 0 now sub 0 in for y in the first equation x + y = 2 x + 0 = 2 x = 2 check... 3x - 2y = 6 3(2) - 2(0) = 6 6 - 0 = 6 6 = 6 (correct) x = 2 and y = 0 (One solution) (b) 3x + y = 1 y = -3x + 1 now sub -3x + 1 in for y in the other equation 2y = 2 - 6x 2(-3x + 1) = 2 - 6x -6x + 2 = 2 - 6x -6x + 6x = 2 - 2 0 = 0 INFINITELY MANY SOLUTIONS (c) 2x - y = 7 2x - 7 = y now sub 2x - 7 in for y in the other equation 4x - 2y = 8 4x - 2(2x - 7) = 8 4x - 4x + 14 = 8 0 = -14 + 8 0 = - 6 NO SOLUTION

OpenStudy (anonymous):

pellet happens! You've done it correctly :( pellet

OpenStudy (texaschic101):

:)

OpenStudy (anonymous):

Can i Have Your Facebook or Twitter? i'll ask you there if i need :) please?

OpenStudy (texaschic101):

I don't have facebook or twitter or even instagram...sorry

OpenStudy (anonymous):

How Can I Contact you if i need an answer? :(

OpenStudy (texaschic101):

if I am on this site when you are, you can tag me

OpenStudy (texaschic101):

or you can fan me and see when I am on this site

OpenStudy (anonymous):

hey?

OpenStudy (anonymous):

I'm Your Fan :)

OpenStudy (texaschic101):

cool :)

OpenStudy (anonymous):

How Can I See You Online? :)

OpenStudy (anonymous):

If I'm Right , You're A Girl? isn't it?

OpenStudy (texaschic101):

I don't really go to social websites...well...except for this one.

OpenStudy (anonymous):

How Can i see you online in this site? :)

OpenStudy (texaschic101):

I am a girl...you are right. You can't SEE me on this site, but you can talk to me.

OpenStudy (anonymous):

Can I ask you a personal question? :)

OpenStudy (texaschic101):

of course

OpenStudy (anonymous):

How Old Are you And What's Your Real name ? :">

OpenStudy (texaschic101):

I do not give out my age. My name is Kellie

OpenStudy (anonymous):

you aged 35 years old or above? :)

OpenStudy (anonymous):

Did I Guess it Right ma'am? :)